# Higher-order homogenised riblet boundary conditions

### v1

@@ 7 Summing-up and numerical examples
$\begin{split}&\left[\begin{matrix}u\\ v\\ w\end{matrix}\right]=\ \varepsilon\left[\begin{matrix}h_{1}u_{z}\\ a_{1}v_{z}\\ 0\end{matrix}\right]+\ \varepsilon^{2}\left[\begin{matrix}{\color[rgb]{.5,.5,.5}h_{2}u_{yz}}+{}\\ {\color[rgb]{.5,.5,.5}a_{2}v_{yz}}+{}\\ {}\end{matrix}\begin{matrix}h_{2}^{(p_{x})}p_{x}+{\color[rgb]{.5,.5,.5}h_{2}^{(v_{xz})}v_{xz}}\\ b_{2}p_{y}+{\color[rgb]{.5,.5,.5}f_{2}u_{xz}}\\ -c_{2}v_{yz}+g_{2}u_{xz}\end{matrix}\right]+{}\\ &\varepsilon^{3}\left[\begin{matrix}h_{3}u_{yyz}+{\color[rgb]{.5,.5,.5}h^{(p_{xy})}_{3}p_{xy}}+h_{3}^{(u_{zt})}(u_{zt}-u_{xxz})+{\color[rgb]{.5,.5,.5}h^{(nl)}_{3}u_{z}v_{z}}+h_{3}^{(v_{xyz})}v_{xyz}+h_{3}^{(u_{xxz})}u_{xxz}\\ a_{3}v_{yyz}+{\color[rgb]{.5,.5,.5}b_{3}p_{yy}}+e_{3}^{(v_{zt})}(v_{zt}-v_{xxz})+{\color[rgb]{.5,.5,.5}e_{3}^{(nl)}v_{z}^{2}}+f_{3}^{(u_{xyz})}u_{xyz}+{\color[rgb]{.5,.5,.5}f_{3}^{(p_{xx})}p_{xx}}+f_{3}^{(v_{xxz})}v_{xxz}\\ -c_{2}v_{yz}+g_{2}u_{xz}+{\color[rgb]{.5,.5,.5}g_{3}^{(u_{xyz})}u_{xyz}}+g_{3}^{(p_{xx})}p_{xx}+{\color[rgb]{.5,.5,.5}g_{3}^{(v_{xxz})}v_{xxz}}\end{matrix}\right]\end{split}$ (7.1)
@@
Shape: triangular rectangular trapezoidal sawtooth slip-no-slip
$h_{1}$ (2.13) 00.170741 00.114319 00.191366 00.127310 00.110325
$a_{1}$ (4.17) 00.080200 00.048000 00.081700 00.073500 00.054700
$h_{2}$ (2.23) 00.000000 00.000000 00.000000 00.000000 00.000000
$h_{2}^{(p_{x})}$ (3.10) -0.028211 -0.021001 -0.034823 -0.014104 00.000000
$h_{2}^{(v_{xz})}$ (6.10) 00.000000 00.000000 00.000000 00.003379 00.000000
$a_{2}$ (4.25) 00.000000 00.000000 00.000000 00.000000 00.000000
$b_{2}$ (4.25) -0.005823 -0.002076 -0.005411 -0.004794 00.000000
$-c_{2}$ (4.19) -0.005823 -0.002076 -0.005411 -0.004794 00.000000
$f_{2}$ (6.13) 00.000000 00.000000 00.000000 -0.003392 00.000000
$g_{2}$ (6.13) -0.028213 -0.021001 -0.034825 -0.014105 00.000000
$h_{3}$ (2.30) 00.001658 00.000497 00.002330 00.000686 -0.000767
$h_{3}^{(u_{zt})}$ (6.15) -0.003593 -0.002356 -0.004819 -0.001505 -0.001050
$h_{3}^{(nl)}$ (6.17) 00.000000 00.000000 00.000000 00.000000 00.000000
$h_{3}^{(v_{xyz})}$ (6.21) -0.002664 -0.001329 -0.003696 -0.000959 -0.000756
$h_{3}^{(p_{xy})}$ (6.21) 00.000000 00.000000 00.000000 -0.000232 00.000000
$h_{3}^{(u_{xxz})}$ (6.21) 00.001898 00.004648 00.002032 00.000986 00.001735
$a_{3}$ (4.32) 00.003013 00.001885 00.002978 00.002348 00.002452
$b_{3}$ (4.32) 00.000000 00.000000 00.000000 -0.000199 00.000000
$-c_{3}$ (4.27) 00.000000 00.000000 00.000000 -0.000200 00.000000
$-d_{3}$ (4.28) 00.001665 00.000981 00.001975 00.000848 00.000000
$e_{3}^{(v_{zt})}$ (5.15) -0.000108 -0.000196 -0.000132 -0.000090 -0.000259
$e_{3}^{(nl)}$ (5.15) 00.000000 00.000000 00.000000 00.000000 00.000000
$f_{3}^{(u_{xyz})}$ (6.24) -0.001350 -0.000559 -0.001833 -0.000551 -0.000751
$f_{3}^{(p_{xx})}$ (6.24) 00.000000 00.000000 00.000000 00.001071 00.000000
$f_{3}^{(v_{xxz})}$ (6.24) -0.000998 -0.000769 -0.001394 -0.001029 -0.000448
$g_{3}^{(u_{xyz})}$ (6.25) 00.000000 00.000000 00.000000 00.000219 00.000000
$g_{3}^{(p_{xx})}$ (6.25) 00.010247 00.009021 00.012171 00.003065 00.000000
$g_{3}^{(v_{xxz})}$ (6.25) 00.000000 00.000000 00.000000 -0.001070 00.000000
Table 1: Up to third-order protrusion heights for some typical riblet shapes.
The coefficients involved in this formula have been computed numerically for five wall shapes, chosen for the sake of example but also for their objective interest:
@@
    Triangular riblets with an equilateral shape, tip angle $60^{\circ}$, and height $\sqrt{3/4}s^{*}$. A relatively common shape in both simulations and experiments.
@@
    Rectangular riblets with height $0.5s^{*}$ and width $0.2s^{*}$. Not particularly effective for drag reduction, but easy to mesh for the purpose of numerical simulation and therefore useful for comparison by other authors.
@@
    Trapezoidal riblets with height $0.5s^{*}$ and tip angle $30^{\circ}$, close to what is empirically considered to provide the best drag reduction with a practical shape.
@@
    Asymmetric riblets with a sawtooth shape, height $0.5s^{*}$, and tip angle such as to yield a width equal to the period $s^{*}$. Devised to test the effects of asymmetry on those coefficients of (7.1) that are mandated to be zero when riblets are left-right symmetric. Some flow in this geometry was computed by Modesti et al. (2021).
@@
    The fifth pattern is not really an indented geometry, but a pattern of slip and no-slip stripes on an impermeable flat surface, with $50\%$$50\%$ surface coverage. This configuration is the simplest idealized model of a superhydrophobic surface when used for drag reduction. It also possesses analytical solutions (Philip, 1972), which provide its two first-order protrusion heights as $h_{1}=\log(2)/(2\pi),a_{1}=h_{1}/2$.
@@
For each of the coefficients in (7.1) a corresponding numerical viscous problem has been solved. This is a less daunting computational task than it may look like because only two solvers have to be programmed, one for a 2D Poisson problem and one for a 2D Stokes problem, with a single set of boundary conditions each, and then applied multiple times with different right-hand sides of either the equations or boundary conditions or both. The equations for either problem have been discretized on a square grid using a second-order, finite-difference, immersed-boundary method (Luchini et al., 2025), enhanced with an analytical correction for the tip singularity. The resulting discrete linear system has been LU-decomposed by direct banded Gauss elimination, just once, and then applied to a battery of different right-hand sides. The values obtained for the various coefficients are collected in table 1.
@@
Figure 2: First-order streamwise velocity (a), spanwise velocity (b), wall-normal velocity (c), and corresponding zeroth-order pressure (d) for the five considered geometries. The riblet profile is drawn on top of each figure.
@@
Figure 3: Second-order contributions to the streamwise velocity. (a) $\overline{\varPhi_{21}}$ of (2.18), (b) $\overline{U_{22}}$ of (3.9), (c) $\overline{U_{23}}$ of (6.8).
@@
Figure 4: Second-order contributions to the spanwise velocity. (a) $\overline{\varPsi_{31}}_{,Z}$ of (4.21), (b) $\overline{\varPsi_{32}}_{,Z}$ of (4.22), (c) $\overline{V_{23}}$ of (6.11).
@@
Figure 5: Third-order velocities ingenerated by the time derivative. (a) streamwise velocity $\overline{U_{33}}$ of (6.14), (b) spanwise velocity $\overline{\varPsi_{43}}_{,Z}$ of (5.12).
@@
Figure 6: Nonlinear third-order velocities. (a) streamwise velocity $\overline{U_{34}}$ of (6.16), (b) spanwise velocity $\overline{\varPsi_{44}}_{,Z}$ of (5.13).
@@
Figure 7: Wall-normal velocity corresponding to the largest wall-normal third-order correction, $\overline{W_{36}}$ defined just below (6.22).
@@
We can remark that, of the 2 first-order, 8 second-order, and 18 third-order potential coefficients, several are zero. For left-right symmetric riblets, which means for the first three columns of table 1, 2 first-order, 4 second-order and 10 third-order coefficients are allowed to be non-zero, as the rest are forbidden by symmetry (as brought out in table 1). In addition, two pairs of second-order coefficients are tied to each other ($b_{2}=-c_{2},h_{2}^{(p_{x})}=g_{2}$), thus making the number of independent second-order coefficients just 2. One can observe in figures 3 and 4 that for each forbidden coefficient the corresponding velocity field is of odd symmetry, thus making its $Y$-average zero.
@@
But even for asymmetric riblets, in the fourth column, astoundingly enough 2 second-order coefficients and 2 third-order coefficients are zero, thus making the totals 2, 6 and 16. This is not just a coincidence or a numerical approximation: as shown in appendix B, these coefficients can actually be proved to be zero for any wall shape. They are also not just any of the terms: $h_{2}$ and $a_{2}$ are the first higher-order protrusion heights that came about in the derivation ((2.23) and (4.25) respectively), and even more importantly $e_{3}^{(nl)}$ and $h_{3}^{(nl)}$ are the only third-order terms that involve nonlinearities of the Navier-Stokes equations ((5.15) and (6.17) respectively). Therefore nonlinearities, which were already barred by symmetry for left-right symmetric wall textures, are actually absent, up to third order, for any texture.
@@
One can also notice from the table that not all coefficients are independent, since $b_{2}=-c_{2}$, $b_{3}=-c_{3}$, $f_{2}=-h_{2}^{(v_{xz})}$, $g_{2}=h_{2}^{(p_{x})}$, $h_{3}^{(p_{xy})}=-g_{3}^{(u_{xyz})}$, and $f_{3}^{(p_{xx})}=-g_{3}^{(v_{xxz})}$. Theoretical justification for some of these symmetries is provided in appendix A.

### v2

@@ 7 Summing-up and numerical examples
$\begin{split}&\left[\begin{matrix}u\\ v\\ w\end{matrix}\right]=\ \varepsilon\left[\begin{matrix}h_{1}u_{z}\\ a_{1}v_{z}\\ 0\end{matrix}\right]+\ \varepsilon^{2}\left[\begin{matrix}{\color[rgb]{.5,.5,.5}h_{2}u_{yz}}+{}\\ {\color[rgb]{.5,.5,.5}a_{2}v_{yz}}+{}\\ {}\end{matrix}\begin{matrix}h_{2}^{(p_{x})}p_{x}+{\color[rgb]{.5,.5,.5}h_{2}^{(v_{xz})}v_{xz}}\\ b_{2}p_{y}+{\color[rgb]{.5,.5,.5}f_{2}u_{xz}}\\ -c_{2}v_{yz}+g_{2}u_{xz}\end{matrix}\right]+\varepsilon^{3}\left[\begin{matrix}h_{3}u_{yyz}+{}\\ a_{3}v_{yyz}+{}\\ {}\end{matrix}\right.\\ &\left.\begin{matrix}{\color[rgb]{.5,.5,.5}h^{(p_{xy})}_{3}p_{xy}}+h_{3}^{(u_{zt})}(u_{zt}-u_{xxz})+{\color[rgb]{.5,.5,.5}h^{(nl)}_{3}u_{z}v_{z}}+h_{3}^{(v_{xyz})}v_{xyz}+h_{3}^{(u_{xxz})}u_{xxz}\\ {\color[rgb]{.5,.5,.5}b_{3}p_{yy}}+e_{3}^{(v_{zt})}(v_{zt}-v_{xxz})+{\color[rgb]{.5,.5,.5}e_{3}^{(nl)}v_{z}^{2}}+f_{3}^{(u_{xyz})}u_{xyz}+{\color[rgb]{.5,.5,.5}f_{3}^{(p_{xx})}p_{xx}}+f_{3}^{(v_{xxz})}v_{xxz}\\ {\color[rgb]{.5,.5,.5}-c_{3}v_{yyz}}-d_{3}p_{yy}+{\color[rgb]{.5,.5,.5}g_{3}^{(u_{xyz})}u_{xyz}}+g_{3}^{(p_{xx})}p_{xx}+{\color[rgb]{.5,.5,.5}g_{3}^{(v_{xxz})}v_{xxz}}\end{matrix}\right]\end{split}$ (7.1)
@@
Riblet shape: $60^{\circ}$-triangular $90^{\circ}$-triangular rectangular $30^{\circ}$-trapezoidal sawtooth slip-no-slip
$h_{1}$ (2.13) 00.170741 00.139685 00.114319 00.191366 00.127310 00.110325
$a_{1}$ (4.18) 00.080237 00.077773 00.048021 00.081519 00.073446 00.055162
$h_{2}$ (2.23) 00.000000 00.000000 00.000000 00.000000 00.000000 00.000000
$h_{2}^{(p_{x})}$ (3.10) -0.028213 -0.016816 -0.021010 -0.034825 -0.014106 00.000000
$h_{2}^{(v_{xz})}$ (6.10) 00.000000 00.000000 00.000000 00.000000 00.003393 00.000000
$a_{2}$ (4.26) 00.000000 00.000000 00.000000 00.000000 00.000000 00.000000
$b_{2}$ (4.26) -0.005822 -0.005800 -0.002077 -0.005398 -0.004789 00.000000
$c_{2}$ (4.20) 00.005820 00.005798 00.002074 00.005396 00.004787 00.000000
$f_{2}$ (6.13) 00.000000 00.000000 00.000000 00.000000 -0.003393 00.000000
$g_{2}$ (6.13) -0.028211 -0.016814 -0.021008 -0.034823 -0.014103 00.000000
$h_{3}$ (2.30) 00.001658 00.000907 00.000497 00.002334 00.000686 -0.000767
$h_{3}^{(u_{zt})}$ (6.15) -0.003593 -0.001960 -0.002356 -0.004819 -0.001505 -0.001050
$h_{3}^{(nl)}$ (6.17) 00.000000 00.000000 00.000000 00.000000 00.000000 00.000000
$h_{3}^{(v_{xyz})}$ (6.21) -0.002620 -0.001222 -0.001331 -0.003638 -0.000977 -0.000758
$h_{3}^{(p_{xy})}$ (6.21) 00.000000 00.000000 00.000000 00.000000 -0.000213 00.000000
$h_{3}^{(u_{xxz})}$ (6.21) 00.001897 00.001015 00.004648 00.002038 00.000994 00.001729
$a_{3}$ (4.33) -0.000589 -0.000420 -0.000131 -0.000628 -0.000412 -0.000377
$b_{3}$ (4.33) 00.000000 00.000000 00.000000 00.000000 00.000196 00.000000
$c_{3}$ (4.28) 00.000000 00.000000 00.000000 00.000000 00.000196 00.000000
$d_{3}$ (4.29) -0.001664 -0.001140 -0.000981 -0.001974 -0.000836 00.000000
$e_{3}^{(v_{zt})}$ (5.14) -0.000575 -0.000540 -0.000295 -0.000572 -0.000442 -0.000262
$e_{3}^{(nl)}$ (5.14) 00.000000 00.000000 00.000000 00.000000 00.000000 00.000000
$f_{3}^{(u_{xyz})}$ (6.24) -0.001351 -0.000724 -0.000559 -0.001832 -0.000551 -0.000758
$f_{3}^{(p_{xx})}$ (6.24) 00.000000 00.000000 00.000000 00.000000 00.001071 00.000000
$f_{3}^{(v_{xxz})}$ (6.24) -0.000998 -0.000694 -0.000770 -0.001394 -0.001029 -0.000454
$g_{3}^{(u_{xyz})}$ (6.25) 00.000000 00.000000 00.000000 00.000000 00.000213 00.000000
$g_{3}^{(p_{xx})}$ (6.25) 00.010247 00.003754 00.009021 00.012171 00.003064 00.000000
$g_{3}^{(v_{xxz})}$ (6.25) 00.000000 00.000000 00.000000 00.000000 -0.001071 00.000000
Table 1: Up to third-order protrusion coefficients for some typical riblet shapes.
The coefficients involved in this formula have been computed numerically for six wall shapes, chosen for the sake of example but also for their objective interest:
@@
    Triangular riblets with an equilateral shape, tip angle $60^{\circ}$, and height $\sqrt{3/4}s^{*}$. A relatively common shape in both simulations and experiments (Choi et al., 1993; Bechert et al., 1997; Endrikat et al., 2021; Modesti et al., 2021; Wong et al., 2024), coefficients also computed in table 4 of Ahmed and Bottaro (2025).
@@
    Triangular riblets with tip angle $90^{\circ}$, and height $0.5s^{*}$. These too considered in (Choi et al., 1993; Bechert et al., 1997; Endrikat et al., 2021; Modesti et al., 2021; Wong et al., 2024). 2nd-order and partial 3rd-order coefficients for this geometry were computed by Bottaro and Naqvi (2020).
@@
    Rectangular riblets with height $0.5s^{*}$ and width $0.2s^{*}$. Not particularly effective for drag reduction, but easy to mesh for the purpose of numerical simulation and therefore useful for comparison with other authors (Endrikat et al., 2021; Modesti et al., 2021; Wong et al., 2024), coefficients also computed in table 4 of Ahmed and Bottaro (2025). García-Mayoral and Jiménez (2011b) considered a similar shape but with tip width $0.25s^{*}$.
@@
    Trapezoidal riblets with height $0.5s^{*}$ and tip angle $30^{\circ}$, close to what is empirically considered to provide the best drag reduction with a practical shape (Endrikat et al., 2021; Modesti et al., 2021; Wong et al., 2024), coefficients also computed in table 4 of Ahmed and Bottaro (2025).
@@
    Asymmetric riblets with a sawtooth shape, height $0.5s^{*}$, and tip angle such as to yield a width equal to the period $s^{*}$. Devised to test the effects of asymmetry on those coefficients of (7.1) that are mandated to be zero when riblets are left-right symmetric. Some flow in this geometry was first simulated by Modesti et al. (2021), coefficients also computed in table 4 of Ahmed and Bottaro (2025).
  6. 6.
@@
    The sixth pattern is not really an indented geometry, but a pattern of slip and no-slip stripes on an impermeable flat surface, with $\%5050\%$ surface coverage. This configuration is the simplest idealized model of a superhydrophobic surface when used for drag reduction. It also possesses analytical solutions (Philip, 1972), which provide its two first-order protrusion heights as $h_{1}=\log(2)/(2\pi),a_{1}=h_{1}/2$. These have been computed by Jelly et al. (2014) and Türk et al. (2014).
Of course the calculation can be easily repeated for different-aspect-ratio or more convoluted riblet shapes; for instance, an alternating sequence of small and large riblets is nothing else than a periodic sequence with double the period. All the numerical codes required are available at https://cplcode.net/Applications/article-CPLcodes/horbcs/.
@@
For each of the coefficients in (7.1) a corresponding numerical viscous problem has been solved. This is less daunting a computational task than it may look like, because only two solvers have to be programmed, one for a 2D Poisson problem and one for a 2D Stokes problem, with a single set of boundary conditions each, and then applied multiple times to different right-hand sides of either the equations or boundary conditions or both. The equations for either problem have been discretized on a square grid using a second-order, finite-difference, immersed-boundary method (Luchini et al., 2025), enhanced with an analytical correction for the tip singularity. The resulting discrete linear system has been LU-decomposed by direct banded Gauss elimination, just once, and then applied to a battery of different right-hand sides. The values obtained for the various coefficients are collected in table 1.
@@
Figure 3: (a) First-order streamwise velocity $\overline{U_{11}}=\overline{\varPhi_{11}}$ of (2.12), responsible for the $h_{1}$ longitudinal protrusion height, (b) spanwise velocity $\overline{V_{11}}=\overline{\varPsi_{21}}_{,Z}$ of (4.15), responsible for the $a_{1}$ transverse protrusion height, (c) wall-normal velocity $\overline{W_{11}}=-\overline{\varPsi_{21}}_{,Y}$, and (d) connected zeroth-order pressure $\overline{P_{01}}$ of (6.6), for the six considered geometries. The corresponding riblet profile is drawn on top of each figure.
@@
Figure 4: Second-order contributions to the streamwise velocity. (a) $\overline{\varPhi_{21}}$ of (2.18), producing coefficient $h_{2}$, (b) $\overline{U_{22}}$ of (3.9), producing coefficient $h_{2}^{(p_{x})}$, (c) $\overline{U_{23}}$ of (6.8), producing coefficient $h_{2}^{(v_{xz})}$.
@@
Figure 5: Second-order contributions to the spanwise velocity. (a) $\overline{\varPsi_{31}}_{,Z}$ of (4.22), producing $a_{2}$ and $c_{3}$, (b) $\overline{\varPsi_{32}}_{,Z}$ of (4.23), producing $b_{2}$ and $d_{3}$, (c) $\overline{V_{23}}$ of (6.11), producing $f_{2}$.
@@
Figure 6: Third-order velocities ingenerated by the time derivative. (a) streamwise velocity $\overline{U_{33}}$ of (6.14), producing $h_{3}^{(u_{zt})}$, (b) spanwise velocity $\overline{\varPsi_{43}}_{,Z}$ of (5.12), producing $e_{3}^{(v_{zt})}$.
@@
Figure 7: Nonlinear third-order velocities. (a) streamwise velocity $\overline{U_{34}}$ of (6.16), producing $h_{3}^{(nl)}$, (b) spanwise velocity $\overline{\varPsi_{44}}_{,Z}$ of (5.13), producing $e_{3}^{(nl)}$.
@@
Figure 8: Wall-normal velocity corresponding to the largest wall-normal third-order correction, $\overline{W_{36}}$ defined just below (6.22), producing $g_{3}^{(p_{xx})}$.
@@
We can remark that, of the 2 first-order, 8 second-order, and 18 third-order potential coefficients, several are zero. For left-right symmetric riblets, which means for the first three columns of table 1, 2 first-order, 4 second-order and 10 third-order coefficients are allowed to be non-zero, as the rest are forbidden by symmetry (as brought out in table 1). In addition, two pairs of second-order coefficients are tied to each other ($b_{2}=-c_{2},h_{2}^{(p_{x})}=g_{2}$), thus making the number of independent second-order coefficients just $2$ for symmetric riblets. One can observe in figures 4 and 5 that for each forbidden coefficient the corresponding velocity field is of odd symmetry, thus making its $Y$-average zero.
@@
But even for asymmetric riblets, in the fifth column, astoundingly enough 2 second-order coefficients and 2 third-order coefficients are zero, thus making the non-zero totals 2, 6 and 16. This is not just a coincidence or a numerical approximation: as shown in appendix B, these coefficients can actually be proved to be zero for any wall shape. They are also not just any of the terms: $h_{2}$ and $a_{2}$ are the first higher-order protrusion coefficients that came about in the derivation ((2.23) and (4.26) respectively), and even more importantly $e_{3}^{(nl)}$ and $h_{3}^{(nl)}$ are the only third-order terms that involve nonlinearities of the Navier-Stokes equations ((5.14) and (6.17) respectively). Therefore nonlinearities, which were already barred by symmetry for left-right symmetric wall textures, are actually absent, up to third order, for any textures.
@@
One can also notice from the table that not all coefficients are independent, since $b_{2}=-c_{2}$, $b_{3}=c_{3}$, $f_{2}=-h_{2}^{(v_{xz})}$, $g_{2}=h_{2}^{(p_{x})}$, $h_{3}^{(p_{xy})}=-g_{3}^{(u_{xyz})}$, and $f_{3}^{(p_{xx})}=-g_{3}^{(v_{xxz})}$. Theoretical justification for some of these identities is provided in appendix A.